OpenStudy (anonymous):

What is the axis of symmetry of the graph of y = -3(x + 8)2 + 5? a. x = -3 b. x = 5 c. x = 8 d. x = -8

OpenStudy (owlcoffee):

\[y=-3(x+8)^2+5\] The axis of symmetry is that line that divides a function into two symmetric parts, but in this case, it is a parabola. I know it's a parabola because there is a variable that is squared in the function, in this case being \((x+3)^2\). But, a parabola has it's symmetry axis superposed to that line composed by all the mid-points of the points defined by the parabola. In other words, you can find the symmetry axis by finding the roots, or "zeroes" of the equation and then calculating the mid-point between them. So, we will take in consideration when \(y=0\) , so therefore: \[-3(x+8)^2+5=0\] So, I'll leave to you the solving for "x" part.

OpenStudy (campbell_st):

the vertex form is \[y = a(x -h)^2 + k\] the vertex is (h, k) where h is the line of symmetry and k is the minimum or maximum value of the curve

OpenStudy (anonymous):

None of that helped me to solve it...but thanks

OpenStudy (campbell_st):

the vertex form \[y = a(x - h)^2 + k\] your question y = -3(x + 8)^2 + 5 match them up h = -8, k = 5 so the vertex is at (-8, 5) the x value in the vertex is the equation of the line of symmetry